OpenStudy (anonymous):

Simplify (1-cosΘ)(1+cosΘ) / (1-sinΘ)/(1+sinΘ)?

OpenStudy (freckles):

\[\frac{(1-\cos(\theta))(1+\cos(\theta))}{\frac{1-\sin(\theta)}{1+\sin(\theta)}} \text{ or } \frac{\frac{(1-\cos(\theta))(1+\cos(\theta))}{1-\sin(\theta)}}{1+\sin(\theta)}\]

OpenStudy (jdoe0001):

\(\bf \large \cfrac{\frac{1-cos(\theta)}{1+cos(\theta)}}{\frac{1-sin(\theta)}{1+sin(\theta)}} ?\)

OpenStudy (freckles):

\[\frac{a}{\frac{b}{c}}=a \cdot \frac{c}{b} \text{ if the first one } \\ \frac{\frac{a}{b}}{c}=\frac{a}{b} \cdot \frac{1}{c} \text{ if the second one }\]

OpenStudy (jdoe0001):

or \(\bf \cfrac{(1-cos(\theta)(1+cos(\theta)}{\frac{1-sin(\theta)}{1+sin(\theta)}} ?\)

OpenStudy (anonymous):

a) sin^2Θ b) cos^2Θ c) tan^2Θ d) cosΘ/sinΘ

OpenStudy (anonymous):

@freckles which one do you think it is?

OpenStudy (freckles):

I don't even know what the problem is

OpenStudy (freckles):

joe and I both asked you and you answered neither one of us

OpenStudy (anonymous):

@freckles what does a, b, and c stand for?

OpenStudy (anonymous):

I think it's d...am I right?

OpenStudy (jdoe0001):

\(\bf \cfrac{(1-cos(\theta)(1+cos(\theta)}{\frac{1-sin(\theta)}{1+sin(\theta)}} \) is what you have right?

OpenStudy (jdoe0001):

well... I don' think we even know what the expression is

OpenStudy (jdoe0001):

is it \(\bf \large \cfrac{\frac{1-cos(\theta)}{1+cos(\theta)}}{\frac{1-sin(\theta)}{1+sin(\theta)}} ?\)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@jdoe0001 so is my answer right?

OpenStudy (jdoe0001):

wel... haven't seen any work done though

OpenStudy (jdoe0001):

I don't see it simplyfing per se so.. got any postings of what you did to get that? you can always attach a file using the [Attach File] button, so we can see it

OpenStudy (freckles):

I kinda would like to know what expression you were given because the expression you told joe you were given doesn't match any of the choices given since you haven't been able to tell us what the expression was, a picture of it would be nice

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

never mind, I got it. thank you

OpenStudy (jdoe0001):

from what you showed above, it sounds that what you have is \(\bf \cfrac{(1-cos(\theta)(1+cos(\theta)}{\frac{1-sin(\theta)}{1+sin(\theta)}}\) as opposed to what you said "yes" to

OpenStudy (jdoe0001):

kinda hard to grab a slippery goose =)